### v1

@@ 3.1 Komatsu’s method and properties of the mollifying Function
Following the approach in [18, 1], we define this approximation by mollifying the function $u(x)=|x|^{\alpha-1}$. Specifically, let $\psi_{\delta,\varepsilon}$ be a smooth mollifier and define $u_{\delta,\varepsilon}\coloneqq u\ast\psi_{\delta,\varepsilon}$. The key properties of this function and its derivatives, which are crucial for applying Ito’s formula and controlling the resulting terms, are summarized in the following lemma.
@@
For any $\varepsilon>0$ and $\delta>1$, one can construct a smooth function $\psi_{\delta,\varepsilon}$ such that:
@@
Finally, we will also use the following identity, which describes the action of the generator $L_{\alpha}$ on $u_{\delta,\varepsilon}$. A proof can be found in [1, Lemmas 34 and 35].
@@
Next, we apply Itô’s formula for jump processes (see, e.g., [2, Theorem 4.4.7] to the process $u_{\delta,\varepsilon}(Y_{t})$. Combined with the Lévy-Itô decomposition (see, e.g., [2, Theorem 2.4.16]) and inequality (3.2), this yields the following decomposition:
@@
The process $M^{\delta,\varepsilon}\coloneqq\left(M_{t}^{\delta,\varepsilon}\right)_{0\leq t\leq T}$ is a true martingale. This can be verified by following the argument in [1, subsection 5.1], which involves decomposing the process into its $L^{1}$ and $L^{2}$ martingale parts. This argument relies on the Lipschitz continuity of $u_{\delta,\varepsilon}$ (which follows from the boundedness of its derivative $u^{\prime}_{\delta,\varepsilon}$) and the boundedness of the jump coefficients (see Subsection 5.3). While the boundedness of $u^{\prime}_{\delta,\varepsilon}$ was already sufficient for this purpose in the zero-drift case, our new, sharper estimate (3.3) for $|u^{\prime}_{\delta,\varepsilon}(x)|$ plays a crucial role in the next step of the proof: controlling the drift term $I_{t}^{\delta,\varepsilon,b}$. It is this refined estimate that makes our extension to the non-zero drift case possible.
@@
This term will be controlled later by taking the expectation, which will connect it to the coefficient distance $B$. The second integral, arising from the path deviation $Y_{s}$, is handled by leveraging the Lipschitz continuity of $\tilde{b}$ along with the same derivative estimate. This will produce a term involving $|Y_{s}|$ which will be controlled using Gronwall’s inequality at the final stage of the proof. This leads to the bound:
@@
Having addressed the drift term, we proceed to analyze the jump compensator term, $I_{t}^{\delta,\varepsilon,\sigma}$. The argument to bound this term is identical to that used in the proof of [1, Theorem 31]. This is because the structure of the jump term analysis, which relies on a change of variables and the properties of the mollifier $\psi_{\delta,\varepsilon}$, is not affected by the presence of the drift term. We outline the main steps. First, by applying a change of variables and Lemma 3.3, the term is rewritten using the mollifier $\psi_{\delta,\varepsilon}$:
@@
With all components of the Ito decomposition (3.5) now bounded, we are now ready to combine the estimates for the components of (3.5). First, note that since $\alpha\in(1,2)$ and $\tilde{\eta}\in[1/\alpha,1]$, the exponent $\alpha\tilde{\eta}-1$ is always non-negative. Substituting the bounds for $\left|I_{t}^{\delta,\varepsilon,b}\right|$ (from (3.6), (3.7), (3.8)) and $\left|I_{t}^{\delta,\varepsilon,\sigma}\right|$ (from (3.9)) into the main decomposition (3.5), and then using the initial estimate (3.4), we arrive at an integral inequality for $\left|Y_{t}\right|^{\alpha-1}$. This inequality is recursive in nature, as the term $\int_{0}^{t}f_{\tilde{b}}(s)|Y_{s}|^{\alpha-1}\mathrm{d}s$ (from (3.8)) appears on the right-hand side. We handle this recursive structure by applying Gronwall’s inequality, which yields an explicit bound for $|Y_{t}|^{\alpha-1}$ in terms of the initial difference, , the martingale term, and the integrals of coefficient differences. The next step is to take the expectation. Since $M_{t}^{\delta,\varepsilon}$ is a martingale, its expectation vanishes. Applying Fubini’s theorem to the remaining terms and taking the supremum over $t\in[0,T]$, we obtain the following explicit inequality:
@@ 4 Convergence in probability of the solution paths
([21, Lemma 5.3]) Let $Q$ be a quasi-martingale on $[0,T]$. Then, for any $h>0$:
@@ 5.1 The time-homogeneous case
    The distances between the coefficients are small in the weighted norms:
@@ 5.2 An alternative framework: The supremum norm case
  • The coefficients $b$ and $\sigma$ of the baseline process process $X$ are no longer required to be continuous or bounded; mere measurability is sufficient.
  • The coefficients $\tilde{b}$ and $\tilde{\sigma}$ of the perturbed process process $\tilde{X}$ are no longer required to be bounded.
@@ 5.3 Proof of the martingale property for $M^{\delta,\varepsilon}$
This section is dedicated to proving that the process $M^{\delta,\varepsilon}$, defined in (3.5), is a true martingale. This property is crucial for the main proofs of Theorem 3.1, 4.1 and 5.2 and Corollary 5.1, as it allows us to eliminate the stochastic integral term when taking expectations. The proof, which follows the argument in [1, subsection 5.1], hinges on the properties of the mollified function $u_{\delta,\varepsilon}$ in introduced Lemma 3.2. Specifically, we need to establish that $u_{\delta,\varepsilon}$ is Lipschitz continuous, which is a consequence of its first derivative being bounded. The following lemma summarizes the necessary properties.
@@
The remaining integral, $\int|x-y|^{\alpha-2}dy$, is finite for any $x\in\real$ because the singularity of the integrand at $y=x$ is of order $\alpha-2>-1$, making it integrable. As we saw in the analysis of Case 3 above, this integral attains its maximum for $x$ within or near the interval $\left[\varepsilon\delta^{-1},\varepsilon\right]$, and this maximum value is finite. Therefore, $|u_{\delta,\varepsilon}^{\prime\prime}|$ is uniformly bounded for all $x\in\real$. This completes the proof of Lemma 5.3. We have established that both the first and second derivatives of $u_{\delta,\varepsilon}$ are uniformly bounded, properties that are essential for the arguments that follow. ∎
@@
$\leq\|u^{\prime}_{\delta,\varepsilon}\|_{\infty}^{2}\left(K^{2}t+\int_{0}^{t}g_{\tilde{\sigma}}^{2}(s)\mathrm{d}s\right)\int_{|z|<1}|z|^{2}\nu(dz).$
This is finite because the integral $\int_{|z|<1}|z|^{2}\nu(dz)=\int_{|z|<1}c_{\alpha}|z|^{1-\alpha}dz$ converges for $\alpha<2$. The same conclusion holds under the supremum norm assumption of Theorem 5.2, where $K^{2}t+\int_{0}^{t}g_{\tilde{\sigma}}^{2}(s)\mathrm{d}s$ is replaced by a constant related to $\int_{0}^{t}\left\|\sigma(s,\cdot)-\tilde{\sigma}(s,\cdot)\right\|^{2}_{\infty}\mathrm{d}s$. Therefore, $\left(M_{t}^{\delta,\varepsilon,2}\right)_{0\leq t\leq T}$ is a true $L^{2}$-martingale (see, e.g., [2, Theorem 4.2.3], ).
@@
$\leq\|u_{\delta,\varepsilon}\|_{\infty}\mathbb{E}\left[\int_{0}^{t}\int_{|z|\geq 1}\left|\left(\sigma(X_{s})-\tilde{\sigma}\left(s,\tilde{X}_{s}\right)\right)z\right|\nu(dz)ds\right]$
$\leq C\int_{|z|\geq 1}|z|\nu(dz).$
This final integral is finite because $\int_{|z|\geq 1}|z|\nu(dz)=\int_{|z|\geq 1}c_{\alpha}|z|^{-\alpha}dz$ converges for $\alpha>1$. Thus, $\left(M_{t}^{\delta,\varepsilon,1}\right)_{0\leq t\leq T}$ is an $L^{1}(\Omega)$-martingale (see, e.g., [2, P.231]).
@@ 5.5 Application to time-homogeneous approximations
For the stochastic integral term, we apply a standard moment inequality for integrals with respect to symmetric $\alpha$-stable processes (see, e.g., [10, Lemma 1]). This, combined with the uniform boundedness of $\sigma_{n}$ from Assumption A(iii), yields:
@@
Applying Theorem 4.1, we can find a further subsequence $(n_{k})_{k\in\mathbb{N}}$ such that the following limit ${X^{\infty}=\lim_{k\rightarrow\infty}X^{(n_{k})}}$ exists almost surely. The convergence of this subsequence to a limit process $X^{\infty}$ holds almost surely in the space of càdlàg functions equipped with the supremum norm, which implies the limit path is well-defined for all $t\in[0,T]$. The limit $X^{\infty}$ is càdlàg since $X^{(n_{k})}$ converges in supremum norm by Theorem 4.1 (see [2, P.140]). We prove that the limit $X^{\infty}$ is a solution of SDE (1.1).
@@
Statements and Declarations
The author have no competing interests to declare that are relevant to the content of this article.

### v2

@@ 3.1 Komatsu’s method and properties of the mollifying Function
While pathwise uniqueness for multi-dimensional $\alpha$-stable driven SDEs has been studied (e.g., Tsuchiya [23]), extending our quantitative stability analysis to higher dimensions presents a significant challenge. Tsuchiya’s work also utilizes an approximation of the function $u(\mathbf{x})=|\mathbf{x}|^{\alpha-1}$, but the approximation is constructed via Fourier analysis and Bessel functions, which is very different from our direct mollification approach. A key technical innovation of our paper is the new, sharp derivative estimate for the mollified function (inequality (3.3)). It is not clear whether a similar parameter-dependent estimate, which is crucial for our drift-term analysis and parameter optimization strategy, can be obtained from the Fourier-based approximation method used in the multi-dimensional setting.
Thus, our analysis in this paper is restricted to the one-dimensional case. We follow the approach in [8, 20] and define the approximation by mollifying the function $u(x)=|x|^{\alpha-1}$. Specifically, let $\psi_{\delta,\varepsilon}$ be a smooth mollifier and define $u_{\delta,\varepsilon}\coloneqq u\ast\psi_{\delta,\varepsilon}$. The key properties of this function and its derivatives, which are crucial for applying Ito’s formula and controlling the resulting terms, are summarized in the following lemma.
@@
Let $u(x)=|x|^{\alpha-1}$ and $\alpha\in(1,2)$. For any $\varepsilon>0$ and $\delta>1$, one can construct a smooth function $\psi_{\delta,\varepsilon}$ such that:
@@
Finally, we will also use the following identity, which describes the action of the generator $L_{\alpha}$ on $u_{\delta,\varepsilon}$. A proof can be found in [20, Lemmas 34 and 35].
@@
Next, we apply Itô’s formula for jump processes (see, e.g., [1, Theorem 4.4.7] to the process $u_{\delta,\varepsilon}(Y_{t})$. Combined with the Lévy-Itô decomposition (see, e.g., [1, Theorem 2.4.16]) and inequality (3.2), this yields the following decomposition:
@@
The process $M^{\delta,\varepsilon}\coloneqq\left(M_{t}^{\delta,\varepsilon}\right)_{0\leq t\leq T}$ is a true martingale. This can be verified by following the argument in [20, Subsection 5.1], which involves decomposing the process into its $L^{1}$ and $L^{2}$ martingale parts. This argument relies on the Lipschitz continuity of $u_{\delta,\varepsilon}$ (which follows from the boundedness of its derivative $u^{\prime}_{\delta,\varepsilon}$) and the boundedness of the jump coefficients (see Subsection 5.3). While the boundedness of $u^{\prime}_{\delta,\varepsilon}$ was already sufficient for this purpose in the zero-drift case, our new, sharper estimate (3.3) for $|u^{\prime}_{\delta,\varepsilon}(x)|$ plays a crucial role in the next step of the proof: controlling the drift term $I_{t}^{\delta,\varepsilon,b}$. It is this refined estimate that makes our extension to the non-zero drift case possible.
@@
This term will be controlled later by taking the expectation, which will connect it to the coefficient distance $B$. The second integral, arising from the path deviation $Y_{s}$, is handled by leveraging the Lipschitz continuity of $\tilde{b}$ along with the same derivative estimate. This will produce a term involving $|Y_{s}|$ which will be controlled using Gronwall’s inequality at the final stage of the proof. Without the sharp control provided by inequality (3.3), we would find it difficult to obtain an integrand in this form, which is necessary for the subsequent application of Gronwall’s inequality. This leads to the bound:
@@
Having addressed the drift term, we proceed to analyze the jump compensator term, $I_{t}^{\delta,\varepsilon,\sigma}$. The argument to bound this term is identical to that used in the proof of [20, Theorem 31]. This is because the structure of the jump term analysis, which relies on a change of variables and the properties of the mollifier $\psi_{\delta,\varepsilon}$, is not affected by the presence of the drift term. We outline the main steps. First, by applying a change of variables and Lemma 3.4, the term is rewritten using the mollifier $\psi_{\delta,\varepsilon}$:
@@
With all components of the Ito decomposition (3.5) now bounded, we are now ready to combine the estimates for the components of (3.5). First, note that since $\alpha\in(1,2)$ and $\tilde{\eta}\in[1/\alpha,1]$, the exponent $\alpha\tilde{\eta}-1$ is always non-negative. Substituting the bounds for $\left|I_{t}^{\delta,\varepsilon,b}\right|$ (from (3.6), (3.7), (3.8)) and $\left|I_{t}^{\delta,\varepsilon,\sigma}\right|$ (from (3.9)) into the main decomposition (3.5), and then using the initial estimate (3.4), we arrive at an integral inequality for $\left|Y_{t}\right|^{\alpha-1}$. This inequality is recursive in nature, as the term $\int_{0}^{t}f_{\tilde{b}}(s)|Y_{s}|^{\alpha-1}\mathrm{d}s$ (from (3.8)) appears on the right-hand side. We handle this recursive structure by applying Gronwall’s inequality, which yields an explicit bound for $|Y_{t}|^{\alpha-1}$ in terms of the initial difference, the martingale term, and the integrals of coefficient differences. The next step is to take the expectation. Since $M_{t}^{\delta,\varepsilon}$ is a martingale, its expectation vanishes. Applying Fubini’s theorem to the remaining terms and taking the supremum over $t\in[0,T]$, we obtain the following explicit inequality:
@@ 4 Convergence in probability of the solution paths
([17, Lemma 5.3]) Let $Q$ be a quasi-martingale on $[0,T]$. Then, for any $h>0$:
@@ 5.1 The time-homogeneous case
    The distances between the coefficients are small in the weighted norms defined in (1.3):
@@ 5.2 An alternative framework: The supremum norm case
  • The coefficients $b$ and $\sigma$ of the baseline process $X$ are no longer required to be continuous or bounded; mere measurability is sufficient. The only condition required for these coefficients is that a weak solution of $X$ exists.
  • The coefficients $\tilde{b}$ and $\tilde{\sigma}$ of the perturbed process $\tilde{X}$ are no longer required to be bounded. Note that $\tilde{X}$ is assumed to have a strong solution.
@@ 5.3 Proof of the martingale property for $M^{\delta,\varepsilon}$
This section is dedicated to proving that the process $M^{\delta,\varepsilon}$, defined in (3.5), is a true martingale. This property is crucial for the main proofs of Theorem 3.1, 4.1 and 5.2 and Corollary 5.1, as it allows us to eliminate the stochastic integral term when taking expectations. The proof, which follows the argument in [20, Subsection 5.1], hinges on the properties of the mollified function $u_{\delta,\varepsilon}$ in introduced Lemma 3.3. Specifically, we need to establish that $u_{\delta,\varepsilon}$ is Lipschitz continuous, which is a consequence of its first derivative being bounded. The following lemma summarizes the necessary properties.
@@
The remaining integral, $\int_{\varepsilon\delta^{-1}}^{\varepsilon}|x-y|^{\alpha-2}dy$, is finite for any $x\in\real$ because the singularity of the integrand at $y=x$ is of order $\alpha-2>-1$, making it integrable. As we saw in the analysis of Case 3 above, this integral attains its maximum for $x$ within or near the interval $\left[\varepsilon\delta^{-1},\varepsilon\right]$, and this maximum value is finite. Therefore, $|u_{\delta,\varepsilon}^{\prime\prime}|$ is uniformly bounded for all $x\in\real$. This completes the proof of Lemma 5.3. We have established that both the first and second derivatives of $u_{\delta,\varepsilon}$ are uniformly bounded, properties that are essential for the arguments that follow. ∎
@@
$\leq\|u^{\prime}_{\delta,\varepsilon}\|_{\infty}^{2}\left(K^{2}t+\int_{0}^{t}g_{\tilde{\sigma}}^{2}(s)\mathrm{d}s\right)\int_{|z|<1}|z|^{2}\nu(\mathrm{d}z).$
This is finite because the integral $\int_{|z|<1}|z|^{2}\nu(\mathrm{d}z)=\int_{|z|<1}c_{\alpha}|z|^{1-\alpha}\mathrm{d}z$ converges for $\alpha<2$. The same conclusion holds under the supremum norm assumption of Theorem 5.2, where $K^{2}t+\int_{0}^{t}g_{\tilde{\sigma}}^{2}(s)\mathrm{d}s$ is replaced by a constant related to $\int_{0}^{t}\left\|\sigma(s,\cdot)-\tilde{\sigma}(s,\cdot)\right\|^{2}_{\infty}\mathrm{d}s$. Therefore, $\left(M_{t}^{\delta,\varepsilon,2}\right)_{0\leq t\leq T}$ is a true $L^{2}$-martingale (see, e.g., [1, Theorem 4.2.3], ).
@@
$\leq\|u_{\delta,\varepsilon}\|_{\infty}\mathbb{E}\left[\int_{0}^{t}\int_{|z|\geq 1}\left|\left(\sigma(X_{s})-\tilde{\sigma}\left(s,\tilde{X}_{s}\right)\right)z\right|\nu(\mathrm{d}z)ds\right]$
$\leq C\int_{|z|\geq 1}|z|\nu(\mathrm{d}z).$
This final integral is finite because $\int_{|z|\geq 1}|z|\nu(\mathrm{d}z)=\int_{|z|\geq 1}c_{\alpha}|z|^{-\alpha}\mathrm{d}z$ converges for $\alpha>1$. Thus, $\left(M_{t}^{\delta,\varepsilon,1}\right)_{0\leq t\leq T}$ is an $L^{1}(\Omega)$-martingale (see, e.g., [1, P.231]).
@@ 5.5 Application to time-homogeneous approximations
For the stochastic integral term, we apply a standard moment inequality for integrals with respect to symmetric $\alpha$-stable processes (see, e.g., [7, Lemma 1]). This, combined with the uniform boundedness of $\sigma_{n}$ from Assumption A(iii), yields:
@@
Applying Theorem 4.1, we can find a further subsequence $(n_{k})_{k\in\mathbb{N}}$ such that the following limit ${X^{\infty}=\lim_{k\rightarrow\infty}X^{(n_{k})}}$ exists almost surely. The convergence of this subsequence to a limit process $X^{\infty}$ holds almost surely in the space of càdlàg functions equipped with the supremum norm, which implies the limit path is well-defined for all $t\in[0,T]$. The limit $X^{\infty}$ is càdlàg since $X^{(n_{k})}$ converges in supremum norm by Theorem 4.1 (see [1, P.140]). We prove that the limit $X^{\infty}$ is a solution of SDE (1.1).